the mean of a binomial distribution is 8 and variance is 3.4. The probability of success is


1
Expert's answer
2021-03-02T05:49:02-0500

"Probability = 1-\frac{variance}{mean}"


"P(success)= 1-\frac{3.4}{8}"


"P(success)=1- 0.425"


"P(success)= 0.575"